OpenStudy (anonymous):

@Sarkar (a^4+4b^4) divided by (a^2-2ab + 2b^2)

OpenStudy (anonymous):

Do you know \( (a^4+4b^4) = \left(a^2-2 a b+2 b^2\right) \left(a^2+2 a b+2 b^2\right) \) ?

OpenStudy (anonymous):

Ah!!thats the key i was missing..

OpenStudy (anonymous):

Oh I got it guys! Thanks so much!!